Two planes are flying around the world. The first plane is flying at 300 mph and is 350 miles ahead of the second plane. The second plane is flying at 400 mph. How long will it take for the second plane to catch the first plane.
Answer by alicealc(293)

distance of the first plane = sA
distance of the second plane = sB
time for both plane = t

distance = speed*time

sB = sA + 350
400t = 300t + 350
400t - 300t = 350
100t = 350
t = 350/10 = 3.5

so the second plane will take 3.5 hours to catch the first plane